OpenStudy (anonymous):

Please help What is the value of x? x+15=8 x=?

OpenStudy (anonymous):

OK, simple trick I use is... 15-8=?

OpenStudy (anonymous):

Add a negative to your answer, get it?

OpenStudy (anonymous):

@WhiteTiger14

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Subtract 15 from the left side and 15 from the right side because of the Addition property of equality. x + 15 =8 -15 -15 x= -7

OpenStudy (anonymous):

@WhiteTiger14

OpenStudy (anonymous):

@madmik I was trying to avoid giving them the answer. ;)

OpenStudy (anonymous):

is it -7 or 7

OpenStudy (anonymous):

Subtraction my bad

OpenStudy (anonymous):

-7.

OpenStudy (anonymous):

@CrazyCountryGirl Lol oops. I was just trying to help :) But it is the subtraction property of equality

OpenStudy (anonymous):

Here's another one, can you work it out? 9-13=

OpenStudy (anonymous):

4

OpenStudy (anonymous):

You got it partially right. It's -4. Do you know why?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Do you know what a negative number is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OK, so what are you have a problem with>